OpenStudy (anonymous):

h(t)=2*3^(t+3); find h(4*2)

OpenStudy (anonymous):

\[h(4*2)=2*3^((4*2)+3)\] \[h(8)=2*3^((8)+3)\] \[h(8)=2*3^(11)\] \[h(8)=2*177147\] \[h(8)=354294\]

OpenStudy (anonymous):

I believe this is correct

OpenStudy (anonymous):

THANK YOU SOO MUCH

OpenStudy (anonymous):

no problem, and if you have any more problems like this, just message me, I don't always keep chat open